### v2

# Stability Regularized Cross-Validation

Abstract
We revisit the problem of ensuring strong test set performance via cross-validation, and propose a nested k-fold cross-validation scheme that selects hyperparameters by minimizing a weighted sum of the usual cross-validation metric and an empirical model-stability measure. The weight on the stability term is itself chosen via a nested cross-validation procedure. This reduces the risk of strong validation set performance and poor test set performance due to instability. We benchmark our procedure on a suite of $13$ real-world datasets, and find that, compared to $k$-fold cross-validation over the same hyperparameters, it improves the out-of-sample MSE for sparse ridge regression and CART by $4\%$ and $2\%$ respectively on average, but has no impact on XGBoost. It also reduces the user’s out-of-sample disappointment, sometimes significantly. For instance, for sparse ridge regression, the nested k-fold cross-validation error is on average $0.9\%$ lower than the test set error, while the $k$-fold cross-validation error is $21.8\%$ lower than the test error. Thus, for unstable models such as sparse regression and CART, our approach improves test set performance and reduces out-of-sample disappointment.
Machine Learning, ICML
1 Introduction
A central problem in machine learning involves constructing models that reliably generalize well to unseen data. One of the most popular approaches is cross-validation as introduced by (Stone, 1974; Geisser, 1975), which selects hyperparameters that perform well on a cross-validation set as a proxy for strong test set performance. Moreover, in the statistical learning literature, there is a broad set of conditions under which the cross-validation loss is a good estimator of the test set error (Plutowski et al., 1993; Arlot & Celisse, 2010; Homrighausen & McDonald, 2013; Hastie et al., 2020; Patil et al., 2021), especially when the sample size is large (Kearns & Ron, 1997; Bousquet & Elisseeff, 2002).
However, both theory (e.g. Shao, 1993; Gupta & Rusmevichientong, 2021) and experiments (e.g. Reunanen, 2003; Rao et al., 2008; Stephenson et al., 2021; Bates et al., 2021) have documented an adaptivity gap between validation and test set performance (although not always observed (Roelofs et al., 2019; Bates et al., 2021)), especially in settings with limited data. Concretely, an adaptivity gap arises when the validation error is systematically lower than the test set error of an ML model.
One explanation for adaptivity gaps is as follows: validation sets yield approximately unbiased estimates of test set performance for a fixed combination of hyperparameters. However, validation scores are random variables and subject to some variance. Therefore, the act of optimizing the validation set error risks selecting hyperparameter combinations that disappoint out-of-sample. In the extreme case where the number of hyperparameters is much larger than the number of samples, optimizing the (cross) validation error can be viewed as training on the (cross) validation set. This phenomenon is well documented in different parts of the statistics and optimization literature, where it is variously called “post-decision surprise”, “out-of-sample disappointment”, “researcher degrees of freedom” or “the optimizer’s curse” (Harrison & March, 1984; Ng, 1997; Smith & Winkler, 2006; Simmons et al., 2011; Van Parys et al., 2021). This raises the question: is it possible to improve the performance of (cross) validation by combating the adaptivity gap, possibly by selecting models with a higher validation error and a lower score according to a second metric.
In this work, we propose a strategy for mitigating out-of-sample disappointment and show that it sometimes improves cross-validation’s performance. Specifically, we propose selecting models according to a weighted sum of their cross-validation error and their empirical hypothesis stability (see Section 2.2). This is motivated by the observation that both the cross-validation error and the hypothesis stability appear in generalization bounds for test set error; thus, minimizing the cross-validation error alone may lead to the selection of high-variance models that perform poorly out of sample.
Our main contributions are threefold. First, we extend a generalization bound on the test set error due to (Bousquet & Elisseeff, 2002) from leave-one-out to $k$-fold cross-validation (kCV). This generalization bound takes the form of the cross-validation error plus a term related to a model’s hypothesis stability. Second, motivated by this (often conservative) bound, we propose regularizing cross-validation by selecting models that minimize a weighted sum of a validation metric and the hypothesis stability, rather than the validation score alone, to mitigate out-of-sample disappointment without being overly conservative. Indeed, models with a low cross-validation error that are stable generalize better than models with a low cross-validation error that are unstable. Moreover, to select the weight in this scheme, we embed the entire scheme within a nested cross-validation procedure. Finally, we empirically evaluate our proposal using sparse ridge regression, CART, and XGBoost, and find that it improves the out-of-sample performance of sparse ridge regression and CART by $4\%$ on average, but has no impact on XGBoost, likely because XGBoost is more stable.
1.1 Motivating Example: Poor Performance of Cross-Validation for Sparse Linear Regression
We illustrate the pitfalls of cross-validation in a sparse ridge regression setting, as studied by (Bertsimas & Van Parys, 2020; Hazimeh et al., 2022; Liu et al., 2023). Suppose we wish to recover a $\tau_{\text{true}}=5$ sparse regressor which is generated from a stochastic process according to the following setup (cf. Bertsimas et al., 2020): we fix the number of features $p$, number of data points $n$, correlation parameter $\rho=0.3$ and signal to noise parameter $\nu=1$, and generate $\bm{X},\bm{y}$ according to a data generation procedure standard to the literature and stated in Appendix A for brevity.
Following the standard cross-validation paradigm, we then evaluate the cross-validation error for each $\tau$ and $20$ values of $\gamma$ log-uniformly distributed on $10^{-3},10^{3}]$, using the Generalized Benders Decomposition scheme developed by (Bertsimas & Van Parys, 2020) to solve each MIO to optimality, which is of the form
$\min_{\bm{\beta}\in\mathbb{R}^{p}}\quad$ $\frac{1}{2\gamma}\|\bm{\beta}\|_{2}^{2}+\frac{1}{2}\|\bm{X}\bm{\beta}-\bm{y}\|_{2}^{2}\ \text{s.t.}\ \|\bm{\beta}\|_{0}\leq\tau,$ (1)
and selecting the hyperparameter combination with the lowest leave-one-out cross-validation error.
Figure 1 depicts each hyperparameter combination’s leave-one-out (left) and test (right) error, in an overdetermined setting where $n=50,p=10$ (top) and an underdetermined setting where $n=10,p=50$ (bottom). We generate equivalent plots for five-fold cross-validation in Figure 2 (Appendix B) and obtain similar results. In the overdetermined setting, cross-validation performs well: a model trained by minimizing the LOOCV (resp. five-fold) cross-validation error attains a test error within $0.6\%$ (resp. $1.1\%$) of the (unknowable) test minimum. However, in the underdetermined setting, cross-validation performs poorly: a model trained by minimizing the LOOCV (resp. five-fold) error attains a test set error $16.4\%$ (resp. $31.7\%$) larger than the test set minimum and is seven orders of magnitude larger (resp. one order of magnitude larger) than its LOOCV estimator. This highlights the danger of optimizing the cross-validation error alone, especially in underdetermined settings.
Figure 1: Leave-one-out (LOOCV, left) and test (right) error for varying $\tau$ and $\gamma$, for an overdetermined setting (top, $n=50,p=10$) and an underdetermined setting (bottom, $n=10,p=50$). In the overdetermined setting, LOOCV is a good estimate of the test error for most values of parameters $(\gamma,\tau)$. In contrast, in the underdetermined setting, LOOCV is a poor approximation of the test error, and estimators that minimize LOOCV ($\gamma=10^{3}$, $\tau=10$) significantly disappoint out-of-sample. Our conclusions are identical when using five-fold cross-validation (Appendix B).
1.2 Literature Review
From a statistical learning perspective, there is significant literature on quantifying the out-of-sample performance of models with respect to their training and validation error, originating with the works by (Vapnik, 1999) on VC-dimension and (Bousquet & Elisseeff, 2002) on algorithmic stability theory. As noted, for instance, by Ban & Rudin (2019), algorithmic stability bounds are generally preferable: they are a posteriori bounds with tight constants that depend on only the problem data. In contrast, VC-dimension bounds are a priori bounds that depend on computationally intractable constants. A key conclusion from both streams of work is that more stable models tend to disappoint less out-of-sample. We refer to Hardt et al. (2016); Stephenson et al. (2021); Bates et al. (2021) for reviews of cross-validation and algorithmic stability.
Relatedly, Kale et al. (2011) introduces a notion of mean-square stability, and shows that, for learning algorithms satisfying this stability property, the variance of the $k$-fold cross-validation error can drop dramatically relative to a single hold-out estimate. Similarly, Lim & Yu (2016) proposes Estimation Stability with Cross-Validation (ESCV) for tuning the Lasso, namely, augmenting cross-validation with an estimation-stability criterion that tends to select smaller and more stable sparse models for the Lasso. Further, Meinshausen & Bühlmann (2010) define stability differently from this paper—in terms of the probability that a variable is selected in a statistical model—and apply their definition to variable selection in regression. All three works are complementary: rather than using stability to justify CV’s accuracy or to stabilize Lasso model selection, we regularize the CV objective with a hypothesis-stability penalty and tune its weight via nested cross-validation.
More recently, the statistical learning theory literature has been connected to distributionally robust optimization by (Ban & Rudin, 2019; Gupta & Rusmevichientong, 2021; Gupta et al., 2024). Indeed, Ban & Rudin (2019) proposes solving newsvendor problems by designing decision rules that map features to an order quantity and obtain finite-sample guarantees on out-of-sample costs of newsvendor policies. Closer to our work, Gupta & Rusmevichientong (2021) proposes correcting solutions to high-dimensional problems by invoking Stein’s lemma to obtain a Stein’s Unbiased Risk Estimator (SURE) approximation of the out-of-sample disappointment. Moreover, they demonstrate that a naive implementation of leave-one-out cross-validation performs poorly with limited data. Building upon this, Gupta et al. (2024) de-bias a model’s in-sample performance by incorporating a variance gradient correction term derived via sensitivity analysis. Unfortunately, it is unclear how to extend their approach to our setting, as it applies to problems with linear objectives over subsets of $0,1]^{n}$.
2 Stability Regularized Cross-Validation
In this section, we propose techniques for improving the performance of cross-validation. First, we define our notation (Section 2.1) and propose a bound on the test set error of a machine learning model in terms of its $k$-fold error and algorithmic stability, which was originally proven for the special case of leave-one-out cross-validation by (Bousquet & Elisseeff, 2002) (Section 2.2). By leveraging this bound, we propose a technique for improving the performance of cross-validation, namely nested cross-validation with stability regularization (Section 2.3)
2.1 Setup and Notation
We consider a generic $k$-fold cross-validation setting for supervised learning problems with $n$ datapoints, and our notation is standard to the machine learning literature. Concretely, we have features $(\bm{x}_{1},\ldots,\bm{x}_{n})\in\mathcal{X}^{n}\subseteq\mathbb{R}^{n\times p}$ and response $(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}\subseteq\mathbb{R}^{n}$, and we assume that the data $(\bm{x}_{i},y_{i})_{i\in[n]}\in\mathcal{X}\times\mathcal{Y}$ are drawn i.i.d. from some (unknown) stochastic process. We aim to understand how abstract models $\bm{\beta}:\mathcal{X}\rightarrow\mathcal{Y}$ trained on the full dataset $(\bm{X},\bm{y})$ generalize to other observations from the same stochastic process, by studying $\bm{\beta}$ and related models $\bm{\beta}^{(\mathcal{N}_{j})}:\mathcal{X}\rightarrow\mathcal{Y}$ trained on all data apart from the points $i\in\mathcal{N}_{j}$. We formalize the notion of generalization with a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, e.g., $\ell(y,\hat{y})=(\hat{y}-y)^{2}$. We let $\{\mathcal{N}_{j}\}_{j\in[k]}$ be a partition of the integers $[n]$ into $k$ disjoint subsets, where $k$ is frequently either $5,10$ (five-fold and ten-fold cross-validation) or $n$ (leave-one-out cross-validation), and the cardinality of each fold $\mathcal{N}_{j}$ is typically identical.
To make the dependence of our abstract models $\bm{\beta}$ on hyperparameters explicit, we let $\bm{\theta}\in\Theta$ be the vector of all hyperparameters that are used to select $\bm{\beta}$, and $\Theta$ denote the set of all possible hyperparameter combinations. We define an abstract model trained using hyperparameters $\bm{\theta}$ by $\bm{\beta}(\bm{\theta})$ for concreteness and assume that this choice is unique for simplicity. Finally, at the risk of overloading notation, we let $\bm{\beta}(\bm{\theta},\bm{x}_{i})\in\mathcal{Y}$ denote the output predicted by model $\bm{\beta}(\bm{\theta})$ with data $\bm{x}_{i}$.
Given the above notation, the $k$-fold cross-validation error with hyperparameters $\bm{\theta}$ is given by:
$h(\bm{\theta})=\frac{1}{n}\sum_{j=1}^{k}\sum_{i\in\mathcal{N}_{j}}\ell\left(y_{i},\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x}_{i})\right)$ (2)
Moreover, for each $j\in[k]$, we let the $j$th partial $k$-fold error be:
$h_{j}(\bm{\theta})=\sum_{i\in\mathcal{N}_{j}}\ell\left(y_{i},\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x}_{i})\right).$ (3)
Therefore, the average $k$-fold error is given by $1/n\sum_{j=1}^{k}h_{j}(\bm{\theta})=h(\bm{\theta})$.
We now define the stability of our models, following (Bousquet & Elisseeff, 2002). Specifically, let $\mu_{h}$ be the hypothesis stability of our learner analogously to (Bousquet & Elisseeff, 2002, Definition 3) but where $k<n$ folds are possible:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mu_{h}:=\max_{j\in[k]}\mathbb{E}_{\mathcal{S}\sim\mathcal{D}}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}$ $\bigg|\ell\left(y,\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x})\right)$ (4)
$-\ell\bigg(y,\beta(\bm{\theta},\bm{x})\bigg)\bigg|,$
i.e., the expectation over the training set $\mathcal{S}$ sampled from an underlying stochastic process $\mathcal{D}$, and any relevant fold split randomness, evaluated at a point $(\bm{x},y)\sim\mathcal{D}$ not necessarily contained in the training set. This quantity measures the worst-case average absolute change in the loss after omitting a fold of data.
Unfortunately, the hypothesis stability involves computing a possibly high-dimensional expectation and thus is #\#P-hard to compute in general (even when $\bm{\beta}(\bm{\theta})$ is computable in polynomial time) by reduction from two-stage stochastic programming with random recourse (Hanasusanto et al., 2016; Bertsimas & Sturt, 2020). Moreover, the stochastic process from which $(\bm{x}_{i},y_{i})$ is drawn is often unknown in practice, and thus cannot be computed even with the help of Monte-Carlo simulation or similar techniques to evaluate high-dimensional integrals. We have the following result:
Proposition 2.1.
Determining the quantity $\mu_{h}$ in (4) is #\#P-hard, i.e., at least as hard as computing the number of optimal solutions to an NP-complete problem, even for the simple case of a binary loss function and i.i.d. discrete random data $(\bm{x}_{i},y_{i})$.
Proof.
We perform a reduction from Bertsimas & Sturt (2020, Corollary 2.1), where the authors report that if $X_{1},\ldots,X_{m}$ are i.i.d. discrete random variables with support containing at least $m$ distinct values, then computing $\mathbb{P}(\sum_{i}X_{i}\leq\alpha)$ is #\#P-hard.
[…]
$(A(T))(x)\;:=\;\begin{cases}\mathbf{1}\{\sum_{z\in T}x(z)\leq\alpha\},&|T|=m,\\ 0,&|T|=2m.\end{cases}$
This algorithm runs in time $O(|T|)$.
Now draw a sample $S=(z_{1},\ldots,z_{2m})\sim\mathcal{D}^{2m}$ and let $S^{(-j)}:=(z_{i})_{i\notin N_{j}}$ denote the sample with fold $j$ removed. By construction, $A(S)=0$, while for each $j\in\{1,2\}$, $A(S^{(-j)})=\mathbf{1}\{\sum_{i\notin N_{j}}x_{i}\leq\alpha\}$, where $x_{i}$ is the feature in $z_{i}$. Since the predictors are constant, the inner expectation over an independent $z\sim\mathcal{D}$ is immaterial and
$\mu_{h}=$ $\max_{j\in\{1,2\}}\mathbb{E}_{S\sim\mathcal{D}^{2m}}\mathbb{E}_{z\sim\mathcal{D}}\Bigl[\bigl|\ell(z,A(S^{(-j)}))-\ell(z,A(S))\bigr|\Bigr]$
$=\mathbb{P}\!\left(\sum_{i=1}^{m}X_{i}\leq\alpha\right)=p.\qed$
Thus, to avoid the cost of computing the potentially unknowable and #\#P-hard quantity $\mu_{h}$, we propose approximating (4) via the empirical hypothesis stability analogously to (Bousquet & Elisseeff, 2002, Definition 3) but where $k<n$ folds are possible:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{h}:=\max_{j\in[k]}\frac{1}{n}\sum_{i=1}^{n}\left|\ell\left(y_{i},\beta^{(\mathcal{N}_{j})}(\bm{\theta},\bm{x}_{i})\right)-\ell\left(y_{i},\beta(\bm{\theta},\bm{x}_{i})\right)\right|.$ (5)
This can be viewed as a sample-average approximation of the hypothesis stability (see Shapiro et al., 2021, for a general theory), and thus it converges almost surely to the true hypothesis stability as $n\rightarrow\infty$ under the usual assumptions of the sample-average-approximation method (Shapiro et al., 2021).
We remark that the above empirical hypothesis stability differs from the pointwise hypothesis stability defined by (Bousquet & Elisseeff, 2002, Definition 4) as we average over all data points, rather than only data points omitted when training $\bm{\beta}^{(\mathcal{N}_{j})}$. This is because the pointwise stability only gives generalization bounds on the training error (Bousquet & Elisseeff, 2002), while we are interested in bounds on the kCV error.
2.2 Generalization Bound
By combining our definitions and notation, letting $M$ represent an upper bound on the loss $\ell(y_{i},\bm{\beta}(\bm{\theta},\bm{x}_{i}))$ for any model $\bm{\beta}(\bm{\theta})$ and any data point $(\bm{x}_{i},y_{i})$ (e.g., if $(\bm{x}_{i},y_{i})$ are drawn from a bounded domain), the following result follows from Chebyshev’s inequality and upper bounds the test error $\mathbb{E}_{(x,y)\sim D}[\ell\left(y,\beta(\bm{\theta},\bm{x})\right)]$ in terms of the $k$-fold cross-validation error and the hypothesis stability (proof deferred to Section B.1):
Theorem 2.2.
Suppose the training data $(\bm{x}_{i},y_{i})_{i\in[n]}$ are drawn from an unknown distribution $\mathcal{D}$ such that $M$ and $\mu_{h}$ are finite constants. Further, suppose $n$ is exactly divisible by $k$ and each $\mathcal{N}_{j}$ is of cardinality $n/k$. Then, the following bound on the test error holds with probability at least $1-\delta$ (over the draw of the training sample, and any fold-partition randomness):
$\mathbb{E}_{(x,y)\sim D}[\ell\left(y,\beta(\bm{\theta},\bm{x})\right)]$ (6)
$\leq\frac{1}{n}\sum_{j\in[k]}h_{j}(\bm{\theta})+\sqrt{\frac{M^{2}+6Mk\mu_{h}}{2k\delta}}.$
Remark 2.3 (To Train or to Validate in (6)).
A similar bound to (6) can be derived using the empirical risk instead of kCV (Bousquet & Elisseeff, 2002, Theorem 11). However, this bound has a larger constant ($12$ instead of $6$) and still involves the expensive pointwise hypothesis stability defined by Bousquet & Elisseeff (2002).
Theorem 2.2 reveals that, if the number of folds $k$ increases with $n$, $M$ is finite, and the hypothesis stability $\mu_{h}$ decreases with $n$, then the kCV error generalizes to the test set with high probability as $n$ becomes large. Moreover, when models have the same cross-validation error, hypothesis stability, and loss bound $M$, training on more folds results in a stronger generalization bound.
We remark that Theorem 2.2 implicitly justifies the use of regularization in machine learning, because regularization implicitly controls the hypothesis stability $\mu_{h}$, leading to better generalization properties when $\mu_{h}$ is lower. Indeed, Bousquet & Elisseeff (2002, Theorem 22) provides a result formalizing this notion in the context of supervised learning for Reproducing Kernel Hilbert Spaces.
Unfortunately, in preliminary experiments, we found that Equation (6)’s bound is often excessively conservative in practice, especially when $n\gg p$. This conservatism stems from using Chebyshev’s inequality in the proof of Theorem 2.2, which is known to be tight for discrete measures but excessively conservative over continuous measures (Bertsimas & Popescu, 2005), especially unimodal continuous measures (Van Parys et al., 2016). Thus, motivated by robust optimization, where probabilistic guarantees are used to motivate uncertainty sets but less stringent guarantees are used in practice to avoid excessive conservatism (see Gorissen et al., 2015, Section 3, for a discussion), we leverage Theorem 2.2 to propose a new approach to cross-validation in the rest of this section.
2.3 Stability Regularization and Nesting
Motivated by the idea that more stable models are less likely to disappoint out-of-sample (Theorem 2.2), we propose selecting models that minimize a weighted sum of the kCV error and the empirical hypothesis stability. This corresponds to selecting $\bm{\theta}$ through the optimization problem
$\bm{\theta}\in\operatorname*{arg\,min}_{\bm{\theta}\in\bm{\Theta}}\frac{1}{n}\sum_{j\in[k]}h_{j}(\bm{\theta})+\lambda\hat{\mu_{h}}(\bm{\theta}).$ (7)
Notably, this procedure still requires a hyperparameter $\lambda$, which the user must select. Accordingly, we invoke nested cross-validation, which has been empirically shown to be significantly less vulnerable to out-of-sample disappointment than regular cross-validation (Cawley & Talbot, 2010; Bates et al., 2021). We perform an outer loop over candidate $\lambda$ values. For each $\lambda$, we run an inner (nested) k-fold cross-validation: on each fold, train a model with that $\lambda$ (using the remaining folds with stability regularization) and measure its validation error. We then choose the $\lambda$ that yields the lowest average outer validation error across the $k$ folds.
Specifically, let $\mathcal{N}_{j,l}:=\mathcal{N}_{j}\cup\mathcal{N}_{l}$ denote the data contained in the $j$th or $l$th fold of the training set. Then, we select $\lambda$ by solving
$\lambda\in$ $\arg\min_{\lambda\in\bm{\Lambda}}\sum_{j\in[k]}\sum_{i\in\mathcal{N}_{j}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{j})}(\bm{\theta}_{j}^{\star},\bm{x}_{i})),$ (8)
where $\bm{\beta}^{(\mathcal{N}_{j})}(\bm{\theta}_{j}^{\star})$ denotes a model trained on all data but the fold $\mathcal{N}_{j}$ with hyperparameters $\bm{\theta}_{j}^{\star}$, and $\bm{\theta}_{j}^{\star}$ denotes an optimal solution to the following lower-level problem, which cross-validates $\bm{\theta}$ with the $j$th fold of the data omitted and $\lambda$ fixed, so that $\bm{\theta}^{\star}_{j}$ is in the argmin of
$\min_{\bm{\theta}\in\bm{\Theta}}\frac{1}{n-|\mathcal{N}_{j}|}\sum_{l\in[k]:l\neq j}\sum_{i\in\mathcal{N}_{l}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{j,l})}(\bm{\theta}))+\lambda\hat{\mu}_{h}^{(-j)}(\bm{\theta}),$
where $\hat{\mu}_{h}^{(-j)}(\bm{\theta})$ is calculated without reference to the $j$th fold of the data, $\mathcal{N}_{j}$, in this case. This corresponds to selecting $\lambda$ in a manner that ensures that models $\bm{\theta^{\star}_{j}}$ perform well on average, as in the standard nested cross-validation paradigm.
Finally, once $\lambda=\lambda^{\star}$ is fixed, we select $\bm{\theta}$ by minimizing (7). This selects models that are robust to omitting one fold of the data and tend to perform well out-of-sample, as estimated by nested cross-validation, at the cost of increased hyperparameter selection time. We formalize this procedure in Algorithm 1 (see Appendix C).
Complexity analysis: In terms of the complexity of Algorithm 1, let $|\Theta|$ be the number of hyperparameter candidates, and let $|\Lambda|$ be the number of candidate stability weights. Then, standard $k$-fold cross-validation (kCV) evaluates each hyperparameter combination $\theta$ using $k$ model fits, for a total of $O(k|\Theta|)$ model calls, plus one at the selected model. On the other hand, Algorithm 1 adds an outer $k$-fold loop to tune $\lambda$; for each $\lambda$, each outer fold $t$, and each evaluated $\theta$, the inner loop fits 1+(k1)=k1+(k-1)=k models to compute both the inner CV score and the stability proxy. Thus, a direct implementation of Algorithm 1 requires $O(|\Lambda|k^{2}|\Theta|)$ model fits to select $\lambda^{\star}$, plus $O((k+1)|\Theta|)$ additional fits to select $\theta^{\star}$ at the chosen $\lambda^{\star}$. Relative to standard kCV, this is an overhead of order $|\Lambda|k$ in the number of fitted models. In the special case where the inner search exhaustively evaluates a fixed finite grid $\Theta$, this can be reduced by a factor of $|\Lambda|$ (for a total overhead factor of $k$) by saving all models and computing the optimal $\lambda$ after one pass of the inner loop. However, for adaptive searches (where the set of evaluated $\theta$ depends on $\lambda$, as occurs in our numerics), the $O(|\Lambda|k^{2}|\Theta|)$ bound captures the worst-case fit count.
3 Numerical Experiments
In this section, we evaluate the numerical performance of the nested regularized cross-validation scheme proposed in Section 2. All experiments were implemented in Julia version $1.9$, using Mosek version $11.0$ to solve all conic optimization problems, and conducted on a MacBook Pro with Apple M3 cores and $36$ GB RAM. Where needed, we used the RCall package to access R version $4.2.1$, which calls certain baseline methods implemented in R, via Julia.
3.1 Ridge Regularized Best Subset Selection
We now benchmark our proposed nested cross-validation scheme on sparse regression for a suite of commonly studied real-world datasets. Specifically, we benchmark a cyclic coordinate descent scheme for $\ell_{0}$$\ell_{2}^{2}$ sparse regression, where we repeatedly solve the following lower-level problem [^1] Note that the terms $\beta_{j}^{2}/z_{j}$ implicitly impose the logical constraints $\beta_{j}=0$ if $z_{j}=0$ using the perspective reformulation trick (Günlük & Linderoth, 2010). for different values of the sparsity parameter $\tau$ and the regularization hyperparameter $\gamma$
$\min_{\bm{\beta}\in\mathbb{R}^{p},\bm{z}\in\{0,1\}^{p}}\;$ $\frac{1}{2\gamma}\sum_{j=1}^{p}\frac{\beta_{j}^{2}}{z_{j}}+\frac{1}{2}\|\bm{X}\bm{\beta}-\bm{y}\|_{2}^{2}\ \text{s.t.}\ \sum_{j=1}^{p}z_{j}\leq\tau,$
using the greedy rounding algorithm described by (Bertsimas et al., 2021) to obtain near-optimal solutions in a practically tractable amount of time. In particular, we iteratively minimize the $k$-fold cross-validation error with respect to $\tau$ (with $\gamma$ fixed) and with respect to $\gamma$ (with $\tau$ fixed) until we either cycle or exceed a limit of $10$ iterations. After converging, we fit a model to the full dataset with the hyperparameters ($\tau^{\star},n_{\text{train}}\gamma^{\star}/n)$, where $n_{\text{train}}$ is the number of observations with one fold of the data left out as in (Liu & Dobriban, 2020), to account for the extra fold when fitting a model to the full dataset. Moreover, we perform the same procedure for nested $k$-fold cross-validation with stability regularization. Note that for our cyclic coordinate descent schemes, we set the largest permissible value of $\tau$ such that $\tau\log\tau\leq n$, because Gamarnik & Zadik (2022, Theorem 2.5) demonstrated that, up to constant terms and under certain assumptions on the data generation process, on the order of $\tau\log\tau$ observations are necessary to recover a sparse model with binary coefficients. In preliminary experiments, we relaxed this requirement to $\tau\leq p$ and found that this did not change the optimal value of $\tau$. Moreover, we use a grid size of $20$ values of $\gamma$ log-uniformly distributed over $10^{-3},10^{3}]$, and $10$ values of $\lambda$ log-uniformly distributed over $10^{-4},10^{4}]$.
We compare against the following methods as a benchmark, using in-built functions to approximately minimize the cross-validation loss, and subsequently fit a regression model on the entire dataset with these cross-validated parameters. Note, however, that we are mainly interested in the performance of sparse ridge regression with and without stability regularization. For all methods, we use five folds:
  • The ElasticNet method in the GLMNet package, with grid search on their parameter $\alpha\in\{0,0.1,0.2,\ldots,1\}$.
  • The Minimax Concave Penalty (MCP) and Smoothly Clipped Absolute Deviation Penalty (SCAD) as implemented in the R package ncvreg, using the cv.ncvreg function and default parameters.
  • The L0Learn.cvfit method implemented in the L0Learn R package (cf. Hazimeh & Mazumder, 2020), with a grid of $10$ different values of $\gamma$ and default parameters otherwise.
For each dataset, we repeat the following procedure ten times to reduce the variance of our results: we randomly split the data into $90\%$ training/validation data and $10\%$ testing data, and report the average sparsity of the cross-validated model, the method’s estimate of the MSE (kCV or nested kCV error) and the average test set MSE (using the same splits for all methods to reduce variance). We also report summary statistics in terms of the average percentage improvement of the nested procedure compared to the MSE without nesting, in order that each dataset is weighted equally. For each method, we retrain on the full training/validation dataset after cross-validation.
Table 1 depicts the dimensionality of each dataset, the average $k$-fold cross-validation error (“CV”) or nested $k$-fold cross-validation error (“nCV”), the average test set error (“MSE”), and the sparsity attained by our coordinate descent scheme without any stability regularization or nesting (kCV), our cyclic coordinate descent with nested stability-regularized cross-validation (nested-kCV), and the performance of MCP on each dataset. We remark that the average (geometric mean over the average for each dataset) runtime was $14.8$ seconds for sparse regression, $592$ seconds for sparse regression with nesting, and under $1$ second for all other methods. Table 4 (Appendix E) also shows the performance of other methods from the literature (SCAD, GLMNet, and L0Learn) in the same datasets.
We measure the improvement from nested cross-validation vs. $k$-fold cross-validation by computing the average MSE ratio for each dataset $d$, $r_{d}:=\text{MSE}_{nested,d}/\text{MSE}_{kCV,d}$, then aggregating by taking the geometric mean of $r_{d}$ to compute an average improvement (accounting for the fact that percentage improvement is an asymmetric measure).
Dataset n p kCV nested-kCV MCP
$\tau$ CV MSE $\tau$ nCV MSE $\tau$ CV MSE
Wine 6497 11 9.3 0.544 0.543 9.3 0.545 0.542 11 0.543 0.542
Housing 506 13 11 23.60 23.68 11 23.73 23.70 11 23.79 23.66
AutoMPG 392 25 18.5 8.524 8.608 18 8.647 8.703 18.7 9.051 8.828
Hitters 263 19 11.7 0.076 0.082 14.7 0.080 0.080 11.5 0.077 0.082
Prostate 97 8 5.1 0.529 0.559 4.8 0.571 0.549 7.1 0.570 0.552
Servo 167 19 13.8 0.746 0.771 15.4 0.795 0.715 12 0.752 0.705
Toxicity 38 9 3.8 0.037 0.054 3.8 0.044 0.057 2.7 0.050 0.060
Steam 25 8 2.8 0.404 0.467 2.8 0.565 0.426 2.7 0.511 0.684
Alcohol2 44 21 3.6 0.210 0.472 2.7 0.266 0.229 2.1 0.232 0.273
TopGear 242 373 41.3 0.037 0.050 34.9 0.055 0.062 8.1 0.057 0.066
Bardet 120 200 23.5 0.007 0.010 25.3 0.009 0.010 5.3 0.009 0.011
Vessel 180 486 28.8 0.016 0.027 29.2 0.024 0.027 2.7 0.036 0.036
Riboflavin 71 4088 13.4 0.163 0.299 13.8 0.269 0.297 7.5 0.319 0.229
Table 1: Average performance of methods across a suite of real-world datasets where the ground truth is unknown (and may not be sparse), sorted by how overdetermined the dataset is ($n/p$), and separated into the underdetermined and overdetermined cases.
We observe that across the overdetermined datasets, nested cross-validation improves the out-of-sample MSE of sparse regression by $10.0\%$ on average, while across the underdetermined datasets it worsens the out-of-sample MSE by $5.92\%$ on average, for an overall improvement of $4.85\%$ on average. Notably, the hyperparameters selected by both methods agree $67\%$ of the time, meaning the average improvement conditioned on instances selected when the hyperparameters selected by both methods disagree is above $10\%$. Moreover, the nested kCV error is, on average, $0.9\%$ smaller than the test set error for the stability-regularized method, while the kCV error is on average $21.8\%$ smaller than the test set error for sparse ridge regression. This is especially visible for the most underdetermined datasets (Vessel and Riboflavin), where the kCV error for sparse ridge regression is lower than all other methods, but this does not translate to better out-of-sample performance. In contrast, the nested kCV error is a substantially more accurate estimator of the test set error.
3.2 Tree-Based Methods
We now empirically validate our nested cross-validation scheme on a suite of tree-based methods, for the same datasets as those studied in the previous section. The goal of this section is to establish that stability-regularization also improves the performance of tree-based methods.
We benchmark cyclic coordinate descent for the following two methods, using $10$ values of $\lambda$ log-uniformly distributed over $10^{-4},10^{4}]$ for our nesting procedure:
  • A Julia implementation of CART (Breiman et al., 1984) via the DecisionTree.jl package, where we iteratively optimize the $5$-fold and nested $5$-fold cross-validation error with respect to the tree depth over a grid of the integers $\{1,\ldots,10\}$ and the min_samples parameter over a grid of the integers $\{2,\ldots,10\}$, with the tree depth initially fixed to $5$ for both approaches. Note that the DecisionTree.jl package holds all unspecified parameters to their default values, and thus all remaining CART parameters will take default parameters in this experiment.
  • A Julia implementation of XGBoost (Chen & Guestrin, 2016) via the XGBoost.jl package, where we iteratively optimize the $5$-fold and nested $5$-fold cross-validation error with respect to the max_depth parameter over a grid of the integers $\{1,\ldots,10\}$ and the subsample parameter over the grid $\{0.01,0.02,0.03,\ldots,1.0\}$, with the tree depth initially fixed to $5$ for both approaches.
Table 2 depicts the dimensionality of each dataset, the average $k$-fold cross-validation error (“CV”) or nested $k$-fold cross-validation error (“nCV”), and the average test set error (“MSE”), with and without nested cross-validation. For both methods, the first two columns correspond to $k$-fold cross-validation, and the last two columns to nested stability-regularized cross-validation. Note that we repeat our train/test split procedure 50 times for each dataset and report the average to reduce the variance in our results.
As in the previous experiment, we measure the improvement from nested cross-validation vs. $k$-fold cross-validation by computing the average MSE across each dataset, normalizing by the MSE of $k$-fold cross-validation on that dataset to compute a percentage improvement across each dataset, and then taking the geometric mean across all datasets (to account for the fact that percentage improvement is an asymmetric measure). The average (geometric mean over averages for each dataset) runtime was $8.07$s (resp. $377.1$s) for XGBoost without (with) nesting, and <0.01<0.01s (resp. <0.01<0.01s) for CART with/without nesting.
Dataset n p CART XGBoost
CV MSE nCV MSE CV MSE nCV MSE
Wine 6497 11 0.534 0.531 0.539 0.531 0.422 0.405 0.429 0.405
Housing 506 13 17.385 19.183 18.912 19.040 11.092 12.464 12.155 12.464
AutoMPG 392 25 15.351 15.878 16.55 16.004 10.36 11.872 11.395 11.628
Hitters 263 19 0.050 0.057 0.055 0.058 0.038 0.039 0.042 0.039
Prostate 97 8 0.717 0.788 0.815 0.796 0.581 0.696 0.648 0.679
Servo 167 19 0.520 0.565 0.656 0.565 0.237 0.237 0.351 0.237
Toxicity 38 9 0.081 0.118 0.092 0.114 0.066 0.081 0.078 0.081
Steam 25 8 1.089 1.308 1.47 1.244 0.91 1.064 1.165 1.118
Alcohol2 44 21 0.850 1.110 0.915 1.071 1.123 1.264 1.289 1.253
TopGear 242 373 0.076 0.087 0.086 0.086 0.051 0.052 0.061 0.052
Bardet 120 200 0.015 0.020 0.017 0.020 0.012 0.014 0.013 0.014
Vessel 180 486 0.063 0.149 0.084 0.136 0.111 0.095 0.136 0.095
Riboflavin 71 4088 0.719 0.979 0.817 0.904 0.404 0.463 0.496 0.465
Table 2: Average performance of methods across a suite of real-world datasets where the ground truth is unknown (and may not be sparse), sorted by how overdetermined the dataset is ($n/p$), and separated into the underdetermined and overdetermined cases. For both methods, the first two columns correspond to $k$-fold cross-validation, and the last two columns to nested stability-regularized cross-validation.
We observe that for XGBoost, there is no benefit to the nested cross-validation procedure (average improvement of $0.6\%$), and we select the same hyperparameters whether or not we account for stability via nested cross-validation on $92\%$ of instances. This is likely because XGBoost generates stable models by default (i.e., omitting one fold of the data hardly changes its predictions), so explicitly accounting for model stability does not improve performance.
However, there is a benefit to nested cross-validation for CART: it improves the out-of-sample MSE by $2.2\%$ on average, with a $1.1\%$ average improvement on overdetermined datasets, and a $4.8\%$ average improvement on underdetermined datasets. Because both methods selected the same hyperparameters on $51\%$ of all instances, this corresponds to an average improvement of over $4\%$ on instances where the two methods did not select the same hyperparameters. Moreover, the average test set MSE is within $6.2\%$ of the average nested cross-validation error for CART (and $7.4\%$ for XGBoost), vs. $24.7\%$ inaccurate for $k$-fold cross-validation (and $8.1\%$ for XGBoost), and thus nested cross-validation also reduces out-of-sample disappointment.
4 Conclusion
In this work, we proposed a new approach to hyperparameter selection, namely selecting hyperparameters that minimize a weighted sum of the cross-validation error and the empirical hypothesis stability, with the weight in the weighted sum selected via a nested cross-validation procedure. Across a suite of real-world datasets, our approach improves the out-of-sample MSE by $4\%$ on average for sparse ridge regression and $2\%$ on average for CART, but does not improve XGBoost’s performance. It also dramatically reduces the user’s out-of-sample disappointment.
Future work could involve developing a tighter bound on the test set error than the one derived by (Bousquet & Elisseeff, 2002). It would also be interesting to investigate the performance of our nested cross-validation procedure across a broader range of contexts (e.g., neural networks), and to more extensively quantify the role that algorithmic stability plays in out-of-sample performance.
Impact Statement
This paper presents work aimed at advancing the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Data Generation Process for Heatmaps
Our data generation process for the motivating example follows the data generation process used by (Bertsimas et al., 2020) and is as follows:
  1. 1.
    The rows of the model matrix are generated i.i.d. from a $p$-dimensional multivariate Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Sigma})$, where $\Sigma_{ij}=\rho^{|i-j|}$ for all $i,j\in[p]$.
  2. 2.
    A “ground-truth” vector $\bm{\beta}_{\text{true}}$ is sampled with exactly $\tau_{\text{true}}$ non-zero coefficients. The position of the non-zero entries is randomly chosen from a uniform distribution, and the value of the non-zero entries is either $1$ or 1-1 with equal probability.
  3. 3.
    The response vector is generated as $\bm{y}=\bm{X\beta_{\text{true}}}+\bm{\varepsilon}$, where each $\varepsilon_{i}$ is generated i.i.d. from a scaled normal distribution such that $\sqrt{\nu}=\|\bm{X\beta_{\text{true}}}\|_{2}/\|\bm{\varepsilon}\|_{2}$.
  4. 4.
    We standardize $\bm{X},\bm{y}$ to normalize and center them.
  5. 5.
    We generate a separate test set of $n_{\text{test}}=10,000$ observations drawn from the same underlying stochastic process to measure test set performance, and normalized using the same coefficients as the training set.
Appendix B Heatmaps From Globally Minimizing Five-Fold Cross-Validation Error
We now revisit the problem setting considered in Figure 1, using five-fold rather than leave-one-out cross-validation (Figure 2). Our conclusions remain consistent with Figure 1.
Figure 2: Five-fold (left) and test (right) error for varying $\tau$ and $\gamma$, for the overdetermined setting (top, $n=50,p=10$) and an underdetermined setting (bottom, $n=10,p=50$) considered in Figure 1. In the overdetermined setting, the five-fold error is a good estimate of the test error for most values of parameters $(\gamma,\tau)$. In contrast, in the underdetermined setting, the five-fold error is a poor approximation of the test error, and the estimator that minimizes the five-fold error ($\gamma=6.15$, $\tau=5$) significantly disappoints out-of-sample.
B.1 Proof of Theorem 2.2
Proof.
The result follows analogously to (Bousquet & Elisseeff, 2002, Theorem 11); the main novelty in this proof compared to (Bousquet & Elisseeff, 2002, Theorem 11) is the use of a more general notion of hypothesis stability. In particular, Bousquet & Elisseeff (2002)’s definition is sufficient to derive the result for leave-one-out cross-validation, but not for $k$-fold cross-validation. Accordingly, we first define all the notation needed for the result, then prove the result.
First, suppose that we fix $\theta\in\Theta$ and let $S=(z_{1},\ldots,z_{n})$ with $z_{i}=(x_{i},y_{i})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{D}$ be an observation drawn from the stochastic process denoted by $\mathcal{D}$. Further, let $\{\mathcal{N}_{j}\}_{j=1}^{k}$ be a (possibly random, data-independent) partition of $[n]$ into $k$ disjoint folds of equal size $m:=|\mathcal{N}_{j}|=n/k$. Next, for any dataset $T$, write $\beta_{T}(\theta)$ for the predictor trained on $T$ using hyperparameters $\theta$. In particular, define the full-sample predictor $\beta(\theta):=\beta_{S}(\theta)$ and, for each $j\in[k]$, the predictor trained with fold $j$ removed by
$S^{(-j)}:=(z_{i})_{i\notin\mathcal{N}_{j}},\qquad\beta^{(\mathcal{N}_{j})}(\theta)\;:=\;\beta_{S^{(-j)}}(\theta).$
Further, for a point $z=(x,y)$ and a predictor $f$, write $\ell(z,f):=\ell(y,f(x))$. Next, define the (conditional) population risk of the full-sample predictor and the $k$-fold CV estimate by
$R\;:=\;\mathbb{E}_{z\sim\mathcal{D}}\big[\ell(z,\beta(\theta))\big],\qquad R_{\mathrm{CV}}\;:=\;\frac{1}{n}\sum_{j=1}^{k}\sum_{i\in\mathcal{N}_{j}}\ell\!\left(z_{i},\beta^{(\mathcal{N}_{j})}(\theta)\right).$
Further, let $B_{j}:=(z_{i})_{i\in\mathcal{N}_{j}}\in(\mathcal{X}\times\mathcal{Y})^{m}$ denote fold $j$ viewed as a block, and define the block loss
$L(f,B)\;:=\;\frac{1}{m}\sum_{r=1}^{m}\ell(z_{r},f).$
Then $0\leq L(f,B)\leq M$ by assumption and we can rewrite
$R_{\mathrm{CV}}\;=\;\frac{1}{k}\sum_{j=1}^{k}L\!\left(\beta^{(\mathcal{N}_{j})}(\theta),B_{j}\right),\qquad R\;=\;\mathbb{E}_{B\sim\mathcal{D}^{m}}\big[L(\beta(\theta),B)\big].$
We now control the block stability using $\mu_{h}$. In particular, fix $j\in[k]$ and let $B=(z_{1},\ldots,z_{m})\sim\mathcal{D}^{m}$ be independent of $S$. Then
$\mathbb{E}_{B}\Big[\big|L(\beta^{(\mathcal{N}_{j})}(\theta),B)-L(\beta(\theta),B)\big|\Big]$ $\leq\frac{1}{m}\sum_{r=1}^{m}\mathbb{E}_{z_{r}\sim\mathcal{D}}\Big[\big|\ell(z_{r},\beta^{(\mathcal{N}_{j})}(\theta))-\ell(z_{r},\beta(\theta))\big|\Big]$
$=\mathbb{E}_{z\sim\mathcal{D}}\Big[\big|\ell(z,\beta^{(\mathcal{N}_{j})}(\theta))-\ell(z,\beta(\theta))\big|\Big].$
Taking $\max_{j\in[k]}$ and then $\mathbb{E}_{S}$ shows that the stability parameter of the learning rule when evaluated with the block loss $L$ is at most $\mu_{h}$. Next, the standard second-moment bound for stable algorithms (see, e.g., (Bousquet & Elisseeff, 2002, Lemma 9)) applied to the sample of $k$ blocks $(B_{1},\ldots,B_{k})$ with bounded loss $L$ and stability parameter $\leq\mu_{h}$ yields
$\mathbb{E}\big[(R-R_{\mathrm{CV}})^{2}\big]\ \leq\ \frac{M^{2}}{2k}+3M\mu_{h},$
where the expectation is over the draw of $S$ (and any fold-partition randomness). Moreover, Markov’s inequality gives
$\mathbb{P}\big(R-R_{\mathrm{CV}}\geq\varepsilon\big)\leq\mathbb{P}\big((R-R_{\mathrm{CV}})^{2}\geq\varepsilon^{2}\big)\leq\frac{\mathbb{E}[(R-R_{\mathrm{CV}})^{2}]}{\varepsilon^{2}}.$
Setting $\varepsilon:=\sqrt{\frac{\frac{M^{2}}{2k}+3M\mu_{h}}{\delta}}$ and rearranging terms then yields the claim. ∎
Appendix C Pseudocode for Nested Stability-Regularized Cross-Validation
We present the pseudocode in Algorithm 1. Note that when $\lambda=0$, our selection rule corresponds precisely to standard nested cross-validation, where the nesting step is used to better estimate the out-of-sample performance but the same hyperparameters are selected as in $k$-fold cross-validation. The purpose of the nesting scheme is to select $\lambda$, after which the second half of Algorithm 1 is exactly standard $k$-fold cross-validation.
Implementation note: Algorithm 1 is written as if $\Theta$, the set of all hyperparameter combinations, is enumerated (for the sake of clarity). In this case, the models $\bm{\beta}$ need only be computed once for each $\Theta$ and can be reused across all $\lambda$. However, if $\Theta$ is explored adaptively via coordinate descent (as in our experiments), then the nesting step in Algorithm 1 is not redundant.
Algorithm 1 Stability-regularized nested $k$-fold cross-validation for model selection
1:Input: Dataset $\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]}\in\mathcal{X}\times\mathcal{Y}$; number of folds $k$; learning algorithm $\boldsymbol{\beta}$; loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{+}$; hyper-parameter grid $\bm{\Theta}$; stability-weight grid $\Lambda$.
2:Output: Outer scores $\{m_{\lambda,1},\dots,m_{\lambda,k}\}$; chosen weight $\lambda^{\star}$; chosen hyper-parameters $\bm{\theta}^{\star}$.
3: Randomly partition the dataset into $k$ disjoint folds indexed by $\mathcal{N}_{1},\ldots,\mathcal{N}_{k}$.
4:for $\lambda\in\Lambda$ do
5:  for $t\leftarrow 1$ to $k$ do
6:   $\mathcal{D}_{\text{rest}}\leftarrow\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]\setminus\mathcal{N}_{t}}$
7:   $\triangleright$ inner loop: hyper-parameter search
8:   for all $\bm{\theta}\in\bm{\Theta}$ do
9:    $\mu(\bm{\theta})\leftarrow 0$
10:    $\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta})\leftarrow\textsc{fit\_model}(\mathcal{D}_{\text{rest}},\bm{\theta})$
11:    for $t_{2}\leftarrow 1$ to $k$ do
12:     if $t_{2}\neq t$ then
13:      $\mathcal{D}_{\text{train}}\leftarrow\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]\setminus(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}$
14:      $\mathcal{D}_{\text{val}}\leftarrow\{(\mathbf{x}_{i},y_{i})\}_{i\in\mathcal{N}_{t_{2}}}$
15:      $\boldsymbol{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(\bm{\theta})\leftarrow\textsc{fit\_model}(\mathcal{D}_{\text{train}},\bm{\theta})$
16:      $s_{t_{2}}(\bm{\theta})\leftarrow\dfrac{1}{|\mathcal{N}_{t_{2}}|}\sum_{i\in\mathcal{N}_{t_{2}}}\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(\bm{\theta},\mathbf{x}_{i})\right)$
17:      $\mu(\bm{\theta})\leftarrow\max\!\Biggl(\mu(\bm{\theta}),\;\dfrac{1}{|[n]\setminus\mathcal{N}_{t}|}\sum_{i\in[n]\setminus\mathcal{N}_{t}}\Bigl|\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(\bm{\theta},\mathbf{x}_{i})\right)-\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta},\mathbf{x}_{i})\right)\Bigr|\Biggr)$
18:     end if
19:    end for
20:    $\tilde{s}(\bm{\theta})\leftarrow\dfrac{1}{k-1}\sum_{\begin{subarray}{c}t_{2}=1\\ t_{2}\neq t\end{subarray}}^{k}s_{t_{2}}(\bm{\theta})$
21:    $\triangleright$ mean inner CV score
22:   end for
23:   $\bm{\theta}^{\star}_{\lambda}\leftarrow\arg\min_{\bm{\theta}\in\bm{\Theta}}\Bigl(\tilde{s}(\bm{\theta})+\lambda\,\mu(\bm{\theta})\Bigr)$
24:   $\triangleright$ pick best regularized score
25:   $m_{\lambda,t}\leftarrow\dfrac{1}{|\mathcal{N}_{t}|}\sum_{i\in\mathcal{N}_{t}}\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta}^{\star}_{\lambda},\mathbf{x}_{i})\right)$
26:   $\triangleright$ evaluate on outer fold
27:  end for
28:  $\bar{m}_{\lambda}\leftarrow\dfrac{1}{k}\sum_{t=1}^{k}m_{\lambda,t}$
29:  $\triangleright$ estimated out-of-sample performance at $\lambda$
30:end for
[…]
33:for all $\bm{\theta}\in\bm{\Theta}$ do
34:  $\mu(\bm{\theta})\leftarrow 0$
35:  $\boldsymbol{\beta}(\bm{\theta})\leftarrow\textsc{fit\_model}(\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]},\bm{\theta})$
36:  for $t\leftarrow 1$ to $k$ do
37:   $\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta})\leftarrow\textsc{fit\_model}(\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]\setminus\mathcal{N}_{t}},\bm{\theta})$
38:   $s_{t}(\bm{\theta})\leftarrow\dfrac{1}{|\mathcal{N}_{t}|}\sum_{i\in\mathcal{N}_{t}}\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta},\mathbf{x}_{i})\right)$
39:   $\mu(\bm{\theta})\leftarrow\max\!\Biggl(\mu(\bm{\theta}),\;\dfrac{1}{n}\sum_{i\in[n]}\Bigl|\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta},\mathbf{x}_{i})\right)-\ell\!\left(y_{i},\boldsymbol{\beta}(\bm{\theta},\mathbf{x}_{i})\right)\Bigr|\Biggr)$
40:  end for
41:end for
[…]
43:return $\min_{\lambda\in\Lambda}\bar{m}_{\lambda},\;\lambda^{\star},\;\bm{\theta}^{\star}$
Appendix D Dataset Description
We use a variety of real datasets from the literature in our computational experiments. The information of each dataset is summarized in Table 3. Note that we increased the number of features on selected datasets by including second-order interactions.
Dataset n p Notes Reference
Housing 506 13 (Gómez & Prokopyev, 2021)
Wine 6497 11 (Gómez & Prokopyev, 2021)
AutoMPG 392 25 (Gómez & Prokopyev, 2021)
Hitters 263 19 Removed rows with missing data Kaggle
$\bm{y}=\log(\text{salary})$
Prostate 97 8 R Package ncvreg
Servo 167 19 One-hot encoding of features (Gómez & Prokopyev, 2021)
Toxicity 38 9 (Rousseeuw et al., 2009)
Steam 25 8 (Rousseeuw et al., 2009)
Alcohol2 44 21 2nd order interactions added (Rousseeuw et al., 2009)
TopGear 242 373 (Bottmer et al., 2022)
Bardet 120 200 (Ye et al., 2018)
Vessel 180 486 (Christidis et al., 2020)
Riboflavin 71 4088 R package hdi
Table 3: Real datasets used.
Our sources for these datasets are as follows:
  • Four UCI datasets: AutoMPG, Housing, Servo, and Wine. We obtained these datasets from the online supplement to (Gómez & Prokopyev, 2021).
  • The alcohol dataset distributed via the R package robustbase. Note that we increased the number of features for this dataset by including second-order interactions.
  • The Bardet dataset provided by (Ye et al., 2018).
  • The hitters Kaggle dataset, after preprocessing the dataset to remove rows with any missing entries, and transforming the response by taking log(Salary), as is standard when predicting salaries via regression.
  • The Prostate dataset distributed via the R package ncvreg.
  • The Riboflavin dataset distributed by the R package hdi.
  • The steamUse dataset provided by (Rousseeuw et al., 2009).
  • The topgear dataset provided by (Bottmer et al., 2022).
  • The toxicity dataset provided by (Rousseeuw et al., 2009).
  • The vessel dataset made publicly available by (Christidis et al., 2020).
Appendix E Supplementary Results for Sparse Ridge Regression
We present supplementary information for the sparse ridge regression experiments in Table 4.
Dataset n p SCAD GLMNet L0Learn
$\tau$ CV MSE $\tau$ kCV MSE $\tau$ CV MSE
Wine 6497 11 11 0.543 0.542 11 0.542 0.542 11 0.542 0.542
Housing 506 13 11.2 24.34 23.69 12.1 23.33 23.79 11 23.48 23.61
AutoMPG 392 25 17.3 8.894 9.028 19.6 8.705 8.880 15.7 8.766 8.979
Hitters 263 19 12.1 0.083 0.085 12.2 0.075 0.080 8.9 0.077 0.082
Prostate 97 8 6.7 0.529 0.557 6.9 0.508 0.569 3.2 0.511 0.547
Servo 167 19 12.3 0.707 0.706 15.8 0.676 0.709 11.3 0.687 0.735
Toxicity 38 9 3.5 0.044 0.055 6.1 0.038 0.054 3.2 0.032 0.063
Steam 25 8 2.9 0.485 0.504 4.1 0.450 0.497 4.4 0.493 0.428
Alcohol2 44 21 2.4 0.249 0.271 5.8 0.222 0.240 7.4 0.200 0.287
TopGear 242 373 9.8 0.053 0.063 31.8 0.044 0.048 44.1 0.050 0.060
Bardet 120 200 9.4 0.009 0.009 32.8 0.007 0.009 35.9 0.007 0.009
Vessel 180 486 9.7 0.034 0.037 40.0 0.018 0.021 14.5 0.023 0.025
Riboflavin 71 4088 16.3 0.331 0.358 88.8 0.195 0.276 39.3 0.205 0.352
Table 4: Baseline sparse regression methods, to be compared against cross-validation and nested cross-validation in Table 1.